Part 3: Quotas

Table: Market for Houses

P

QD

QS

$100

1300

100

$200

1100

300

$300

900

500

$400

700

700

$500

500

900

$600

300

1100

$700

100

1300

Notes: P = Price in thousands of $

QD = quantity demanded in thousands of homes

QS = quantity supplied in thousands of homes

  1. Suppose the government implements new zoning regulations that allow a maximum of 300,000 new homes to be constructed. At the quota limit, what is the demand price? What is the supply price? (2 points)
  2. What is the quota rent per home sold? How much total quota rent is extracted by producers? Show these amounts on a graph. (2 points)
  3. Who would gain/lose from this policy? Examine this issue critically by calculating the relevant surplus amounts. (2 points)
  4. Suppose instead, the government revised the zoning standard to allow 500,000 new homes to be constructed. What would happen in the market under this quota? Is this better or worse than the previous rule? Discuss. (2 points)
Answer #1

Figure 1

Market for Houses Quota 1 Quota 2 Price Quota Rent B — - Demand Supply 0 100 200 300 400 500 600 900 1000 1100 1200 1300 1400

At the quota limit of Q=300, the demand price is $600 and the supply price is $200.

Therefore, at this quota limit, the quota rent is ($600-$200)=$400.

The total amount of quota rent that the produce collects is $400×300000=$120000000. This is the blue area between demand and the supply curve up to quota limit Q=300.

From this quota, the producer wins and the consumer loses. and at the end, the society loses as a whole, because quota rent causes deadweight loss to the society. Before the quota, the free market equilibrium price was $400 and the quantity was 700,000. The quota increases the price to $600 and the consumer loses the surplus: (600-400)×300000+0.5×(600-400)×(700000-300000)=$100000000. On the other hand, the producer gains $120000000 and loses 0.5×200×400000=$40000000. Then the total gain to producers is $80000000. The loss to society is $20000000.

If the government revised the standard to 500, then the supply price is $300 and the demand price is $$500. The quota rent falls to $200. This decreases the quota rent to $100000000. This quota is better than the previous one because this decreases deadweight loss by area A in the figure above.